OpenStudy (anonymous):

prove sin^(-1) -x = -sin^(-1) x

OpenStudy (anonymous):

this what i done so far \[\sin ^{-1} -x =m\] -x=sin m x=-sin m

OpenStudy (anonymous):

prove: if \(f\) is odd, then \(f^{-1}\) is odd

Nnesha (nnesha):

sin(-x) = -sin(x) sin and tan are odd function cos (-x) = cos(x) is even

OpenStudy (anonymous):

do it without reference to any property of sine other than that it is odd, meaning \[\sin(-x)=-\sin(x)\]

OpenStudy (anonymous):

thanks people

OpenStudy (loser66):

if sin (a) = x, then sin (-a) = -x, right?

OpenStudy (anonymous):

yup

OpenStudy (loser66):

\(sin (a) = x \implies sin^{-1}(x) = a\) \(sin(-a) = -x\implies sin^{-1}(-x) = -a\), OK?

OpenStudy (anonymous):

ok

OpenStudy (loser66):

Hence from the first guy, you just multiple -1 both sides, you get \(-sin^{-1} (x) = -a\) put it into the second one, you have what you want.

OpenStudy (loser66):